a boat takes 9 hrs to travel from A to B upstream. if river current is 3 km/hr , how long will it take to travel down stream between B and A?
Answer by ankor@dixie-net.com(22740)

boat takes 9 hrs to travel from A to B upstream.
if river current is 3 km/hr, how long will it take to travel down stream between B and A?
:
I think we need to know the A to B distance
Let's assume the distance from A to B = 180 mi
Let s = boat speed in still water
then
(s-3) = effective speed upstream (A to B)
Find s
9(s - 3) = 180
simplify, divide by 9
s - 3 = 20
s = 20 + 3
s = 23 mph speed in still water
:
(s+3) = effective speed downstream (B to A)
Let t = time required to go downstream
t(23 + 3 ) = 180
26t = 180
t = 180/26
t = 6.923 hrs to go downstream (B to A)